OpenStudy (paigejones):

Henry receives a flat rate of $100 per day and can earn a $2 commission for each computer repair (r) he performs. Yolanda receives a flat rate of $80 per day and can earn $5 per computer sale (s) she makes. How would the two equations model the wages Henry and Yolanda make each day?

OpenStudy (yuna1234):

Henry= 100+2r Yolanda= 80+5s